OpenStudy (blondey_315):

Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2). (Points : 1) y = 3x - 5 y = 3x - 2 y = 3x + 1 y = 3x + 3 Question 6. 6. Write an equation in slope-intercept form for the line shown. (Points : 1) y = 2/3 x + 2 y = 3/2 x - 4 y = 3x - 4 y = 2x - 1 Question 7. 7. Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1). (Points : 1) y = -4x + 1 y = -1/2 x + 9 y = 4x + 1 y = -2x

OpenStudy (blondey_315):

Sorry this is so long they are the ones im having trouble with! ho ever helps gets a medal!

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 1}}\quad ,&{\color{blue}{ -2}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies 3 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\) is 6.6 supposed to be the slope-intercept of the one above it?

OpenStudy (blondey_315):

no sorry it was just representing that this is question 6

OpenStudy (jdoe0001):

Write an equation in slope-intercept form for the line shown. <--- where's the "line shown"? then

OpenStudy (blondey_315):

sorry i meant to include that!

OpenStudy (jdoe0001):

6.6 would then be \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 0}}\quad ,&{\color{blue}{ -4}})\quad &({\color{red}{ 2}}\quad ,&{\color{blue}{ -1}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (jdoe0001):

and for 7.7 would be \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ 1}})\quad &({\color{red}{ 5}}\quad ,&{\color{blue}{-1}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (jdoe0001):

notice, all you do is, get the slope plug it in the point-slope form along with the coordinates of either given point(s) and solve for "y"